OpenStudy (anonymous):

Find the y-intercept. -2x - y = -10

hero (hero):

The y-intercept occurs when x = 0. Thus: -2x - y = -10 -2x + 10 = y -2(0) + 10 = y 0 + 10 = y 10 = y